OpenStudy (anonymous):

Write the equation of the line that is tangent to the circle (x + 6)2 + (y + 4)2 = 25 at the point (-9, -8).

OpenStudy (anonymous):

I really would love an explanation, I check google and people get numbers that seem to come out of no where

OpenStudy (anonymous):

There is a neat trick in solving these, write the equation like this \[(x_1+6)(x+6)+(y_1+4)(y+4)=25 \] Now match the x_1, y_1 value with the coordinates of your point and evaluate

OpenStudy (turingtest):

that's a great trick @Spacelimbus never seen it before

OpenStudy (anonymous):

okay, so I get (-3)(x+6) + (-4)(x+4) = 25 \so do I distribute ?

OpenStudy (anonymous):

exactly @Whereisalexis2u

OpenStudy (anonymous):

opps theres a y next to 4

OpenStudy (anonymous):

@TuringTest, it's really cool, I lost the proof long ago though heehe

OpenStudy (anonymous):

all my answer options have 3/4x or 4/3 x

OpenStudy (anonymous):

Hmm @Whereisalexis2u, did I read your equation right though? or did you correct something in the meanwhile? You can check my wolfram graphing link above, it's obviously correct, my solution is: \[ 4y+3x=-59\] so it doesn't go through the origin, but it would be -3/4x

OpenStudy (anonymous):

from there I get 59/3 or 4 ones negative and postive. Youre right however, all the numbers you put down are there

OpenStudy (anonymous):

just in a different way.

OpenStudy (anonymous):

hmm strange *smiles* but okay then (-:

OpenStudy (anonymous):

there should only be one tangent though, not two, two would make no sense.

OpenStudy (anonymous):

it's after all a unique point on the circle, therefore the tangent will be unique.

OpenStudy (anonymous):

no im saying my answer option is between 59/3 or 59/4 59/3 is postive and 59/4 is negative

OpenStudy (anonymous):

oh ok \[ y = -\frac{3}{4}x - \frac{59}{4} \]

OpenStudy (anonymous):

I dont get why its that though.

OpenStudy (anonymous):

this whole sections has been a blur to me

OpenStudy (anonymous):

some are tricks, a lot is geometry, often it really helps if you try to make a picture of it, it's annoying but it helps.

OpenStudy (anonymous):

another way would be implicit differentiation and then derive the slope from that.